Bibliography

[1] Louis Boutet de Monvel, Boundary problems for pseudo-differential operators, Acta Math. 126 (1971), no. 1-2, 11–51.

[2] Gerd Grubb, Complex powers of pseudodifferential boundary value problems with the transmission property, Pseudodifferential operators (Oberwolfach, 1986), Lecture Notes in Math., vol. 1256, Springer, 1987, pp. 169–191.

[3] , Functional calculus of pseudodifferential boundary problems, second ed., Progress in Mathematics, vol. 65, Birkhäuser Boston Inc., Boston, 1996.

[4] R. Mazzeo, The Hodge cohomology of a conformally compact metric, J. Diff. Geom. 28 (1988), 309–339.

[5] R. Mazzeo and R.B. Melrose, Meromorphic extension of the resolvent on complete spaces with asymptotically constant negative curvature, J. Funct. Anal. 75 (1987), 260–310.

[6] , The adiabatic limit, Hodge cohomology and Leray’s spectral sequence, J. Diff. Geom. 31 (1990), 185–213.

[7] R.B. Melrose, Geometric scattering, Cambridge University Press, 1995.